# Gradient-based Sample Selection for Faster Bayesian Optimization

### v1

@@ 3. Overall Regret
Summarizing both phases, the total regret satisfies $R_{N}\;=\;R_{M}^{(\text{full})}\;+\;R_{N-M}^{(\text{selected})}$
@@
$\Pr\left(R_{N}\leq\sqrt{\,C\,M\,\beta_{M}\,\gamma_{M}+2}+\sqrt{\,C\,(N-M)\,\beta_{(N-M)}\,\gamma_{(N-M)}}\quad\forall{N}\geq 1\right)\geq 1-\delta.$
Appendix B Experiments
B.1 GP Study Experiments
Figure 6 and 7 compares the cumulative runtime over $400$ iterations on Eggholder2, Levy20, Rastrigin100 functions and NAS experiment.
Figure 6: Cumulative time cost of algorithms 2 (in seconds).
Figure 7: Cumulative time cost of algorithms 3 (in seconds).
Figure 8 illustrate a GP performance comparison between the GSSBO and the Standard GP-UCB, showing that the quantified differences between the two GPs are minimal. Figure 8 compares the root mean square error (RMSE) in 10,000 samples of the posterior mean functions of the two GPs at each iteration. The observed differences increase sub-linearly, indicating that the disparity between the two methods remains bounded, which is consistent with our theoretical analysis. These results suggest that the GSSBO effectively reduces computational costs while maintaining inference accuracy comparable to, or not significantly lower than, that of the Standard GP-UCB. This outcome validates the efficacy of our proposed approach.
Figure 8: Inference gap between GSSBO and GP-UCB.

### v2

@@ 3. Overall Regret
Summarizing both phases, the total regret satisfies $R_{T_{N}}\;=\;R_{{T_{M}}}^{(\text{full})}\;+\;R_{{T_{N}}-{T_{M}}}^{(\text{selected})}$
@@
$\Pr\left(R_{{T_{N}}}\leq\sqrt{\,C_{1}\,{T_{M}}\,\beta_{T_{M}}\,\gamma_{T_{M}}}+2+\sqrt{\,C_{1}\,({T_{N}}-{T_{M}})\,\beta_{({T_{N}}-{T_{M}})}\,\gamma_{({T_{N}}-{T_{M}})}}\quad\forall{{T_{N}}}\geq 1\right)\geq 1-\delta.$ (58)
Appendix B Appendix Experiments
B.1 Supplementary experiments
Figure 6 and 6 compares the cumulative runtime over $1000$ iterations on Eggholder2, Levy20, Rastrigin100 functions and NAS experiment (in seconds).
Figure 5: Cumulative time cost of algorithms 2.
Figure 6: Cumulative time cost of algorithms 3.
B.2 Subset Samples Distribution Study
Figure 7 illustrates the sample distribution of GSSBO and Standard GP-UCB, on the first two dimensions of the Hartmann6 function. In this experiment, we recorded the first 200 sequential samples from a standard BO process and constrained the buffer size to 100. The objective is to identify the global minimum, and darker-colored samples correspond to values closer to the optimal. During the optimization process, only a small number of samples are located near the optimal value. As shown in the middle panel, the gradient-based sample selection method selects a more informative and diverse subset, retaining more samples closer to the optimal or suboptimal, which is indicated by preserving a higher number of darker-colored samples in the figure. In contrast, the random selection strategy reduces the sample density uniformly across all regions, leading to a significant loss of samples near the optimal or suboptimal, as represented by the retention of many lighter-colored samples in the right panel.
While BO algorithms are theoretically designed to balance exploitation and exploration, with limited budget in practice, they can over-exploit current best regions before shifting to exploration [49], leading to suboptimal performance in locating the global optimums. With gradient-based sample selection, the relative density of samples near the optimal and suboptimal regions increases, maintaining a more balanced distribution. This subset encourages subsequent iterations to focus on regions outside the optimal and suboptimal regions, promoting the exploration of other regions of the search space. As a result, the over-exploitation issue is mitigated.
Figure 7: Sample distribution: GP-UCB (left), GSSBO (middle), and RSSBO (right).
B.3 Experiments on Kmeans++ selection
Oglic [42] and Hayakawa [43] proposed to select a subset in RKHS, then employed them to construct the Nyström low-rank approximation. We included it as an additional baseline, experimental results in Fig.8 demonstrate that K-means++ has no clear advantage in cumulative regret compared to our proposed method in subset selection, and it costs more time than our method.
Figure 8: Cumulative regret of GSSBO and Kmeans++ on the Eggholder2, Hart6, Levy20, Powell50, Rastrigin100 functions and NAS experiment.
B.4 Experiments on high-dimensional BO methods
We noticed that all these algorithms struggled with high-dimensional tasks, so we included some high-dimensional baselines: REMBO [22] and HESBO [23]. The results are shown in Fig. 9, which shows that the difference between the three methods is subtle on Levy20, but REMBO and HSEBO do not perform well on higher-dimensional tasks because these two dimensionality reduction methods do not use all the dimensional information.
Figure 9: Cumulative regret of GSSBO, REMBO, and HSEBO on the Levy20, Powell50, Rastrigin100 functions.
B.5 Experiments on other surrogates
Although our theoretical analysis focuses on Gaussian Process surrogates, our proposed gradient-based sample selection method is not limited to GPs. The core idea of the gradient-based sample selection method is to use gradient information to measure the importance of each sample to model fitting, which is model-independent. Therefore, this method can also be applied to Bayesian neural networks (BNN) [50] and Deep Kernel(DK) methods [51]. We have extended our experiments to other popular surrogate models such as Deep Kernels and Bayesian Neural Networks. Our additional results demonstrate that the proposed sample selection strategy yields performance improvements across these different surrogate models. The updated experimental results can be found in Fig. 11 and Fig. 11.
Figure 10: Cumulative regret of Deep Kernal UCB algorithms on the Hart6 experiment.
Figure 11: Cumulative regret of BNN UCB algorithms on the Hart6 experiment.